OpenStudy (anonymous):

How to count initial value for: f' + 4y = 12e^(2t), y(0) = 5

OpenStudy (turingtest):

you mean y' + 4y = 12e^(2t) ?

OpenStudy (anonymous):

yeah, sorry

OpenStudy (turingtest):

so you have the general solution already?

OpenStudy (anonymous):

well, thats all I know, and I want to know the initial value

OpenStudy (turingtest):

y(0)=5 is what the initial value is here in general, y(0)=whatever is the initial value

OpenStudy (anonymous):

okay, I want to know y(t)

OpenStudy (anonymous):

or how to count it, I know the answer

OpenStudy (turingtest):

that means you want to solve the IVP this is a linear equation, so you can use an integrating factor have you tried that?

OpenStudy (anonymous):

yeah, i want to solve the Initial value problem, but i do not understand what you mean with "you can use an integrating factor"

OpenStudy (turingtest):

do you know any techniques on how to solve linear DE's ? I don't know where to start if you have never heard of an integrating factor perhaps it's just a difference in terminology, I do not know your what native language is does\[\large\mu(t)=e^{\int p(t)dt}\]look familiar as an integrating factor?

OpenStudy (anonymous):

Well my native language is swedish but I study in Finnish :P OK, I think that I should learn that formula; what does it mean?

OpenStudy (turingtest):

this is really best explained through reading up on it here is a good link http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx I suggest you read it, along with perhaps other sections of that site, in order to understand this type of problem

OpenStudy (anonymous):

I tried to search that site for help but did not find a fitting article, Thanks! Will read it.

OpenStudy (turingtest):

if you read that page (which you could not possibly have done in this amount of time) you would see that it shows you exactly how to solve this kind of problem it even derives how it works look, I will work through this one with you, but read the page after!

OpenStudy (turingtest):

one sec...

OpenStudy (turingtest):

in general, say we have the form\[y'+p(t)y=q(t)\]that is called a linear DE the way this is solved is by first multiplying the whole thing through by something called an integrating factor \(\mu(t)\) defined as\[\large\mu(t)=e^{\int p(t)dt}\]still with me?

OpenStudy (anonymous):

I guess..

OpenStudy (turingtest):

well let's try it: your problem is\[y'+4y=12e^{2t}\]so what are \(p(t)\) and \(q(t)\) in this case?

OpenStudy (anonymous):

p(t) is 4 and q(t) is 12e^(2t)?

OpenStudy (turingtest):

exactly :D now since I have given you a formula for the integrating factor \(\mu(t)\) you should be able to tell me what that is too what is it?

OpenStudy (anonymous):

µ(t) = e^(4t)?

OpenStudy (turingtest):

excellent! now multiply the whole equation by that integrating factor and show me what you get (write both sides of the expression please)

OpenStudy (anonymous):

do you mean: e^(4t) * y' + 4ye^(4t) = 12e^(2t) * e^(4t)

OpenStudy (turingtest):

yes, which we can simplify to\[e^{4t}y'+4ye^{4t}=12e^{6t}\]right? now here is here we need to be observant to understand the following step...

OpenStudy (anonymous):

yeah

OpenStudy (turingtest):

look carefully at the left-hand-side notice that it can be written as a result of the product rule\[e^{4t}y'+4ye^{4t}=\frac d{dt}(e^{4t}y)\]so we can rewrite this whole expression as\[\frac d{dt}(e^{4t}y)=12e^{6t}\]this will \(always\) be the case when we do this integrating factor technique!

OpenStudy (turingtest):

(in general we will always get the form\[\frac d{dt}(\mu(t)y)=q(t)\mu(t)\]from which we will now proceed) still with me up to\[\frac d{dt}(e^{4t}y)=12e^{6t}\]?

OpenStudy (anonymous):

this, i did not understand... thinking...

OpenStudy (turingtest):

yes, this is the tricky bit I suppose... -you were with me up to\[e^{4t}y'+4ye^{4t}=e^{6t}\](after multiplying by the integrating factor) right?

OpenStudy (anonymous):

12e^6t, yes

OpenStudy (turingtest):

^right, my bad -and just looking at the LHS, do you not agree that (by what may seem to be a pure coincidence) it is a fact that\[e^{4t}y'+4ye^{4t}=\frac d{dt}(e^{4t}y)\]?

OpenStudy (anonymous):

nooo....?

OpenStudy (turingtest):

product rule:\[(uv)'=u'v+v'u\]let \(u=e^{4t}\) and \(v=y\) \[(e^{4t}y)'=(y)'e^{4t}+y(e^{4t})'=y'e^{4t}+4ye^{4t}\]hence\[y'e^{4t}+4ye^{4t}=(e^{4t}y)'\]

OpenStudy (anonymous):

\[y(e^{4t})' \] be \[4tye^{4t}\] ?

OpenStudy (turingtest):

chain rule:\[(e^{4t})'=(e^{4t})'(4t)'=4e^{4t}\]

OpenStudy (anonymous):

sorry ofcourse :)

OpenStudy (turingtest):

it happens :)

OpenStudy (anonymous):

ok, I'm with you

OpenStudy (turingtest):

so... now you comprehend\[e^{4t}y'+4ye^{4t}=\frac d{dt}(e^{4t}y)\]which setting equal to the RHS again gives\[\frac d{dt}(e^{4t}y)=12e^{6t}\]now this is a separable DE, which you probably have done in calc I

OpenStudy (turingtest):

I think it is best just to refresh your memory and demonstrate the next few steps, so....

OpenStudy (turingtest):

\[\frac d{dt}(e^{4t}y)=12e^{6t}\]\[d(e^{4t}y)=12e^{6t}dt\]\[\int d(e^{4t}y)=12\int e^{6t}dt\]now note the LHS, being the integral of a differential, just becomes the integrand. so we are left with\[e^{4t}y=12\int e^{6t}dt\]so once we can do the integral we are very close to solving for y! why don't you try to do the integral and solve for y yourself?

OpenStudy (anonymous):

e^(4t) * y = 12 * 6 e^(6t) y = 72*e^(2t)

OpenStudy (turingtest):

gotta be way more careful with that calculus\[\int e^{6t}dt\neq6e^{6t}\]and perhaps even more important, you forgot to add the constant +C from the indefinite integral DE's is where that constant you learn to add in calc I really comes into play try that integral again...

OpenStudy (anonymous):

e^(4t) * y = 12 * (1/6) e^(6t) + C y = 2e^(2t) + C*e^(-4t)

OpenStudy (turingtest):

and that is the general solution! nice job :D

OpenStudy (anonymous):

c = 3

OpenStudy (anonymous):

y(t) = 2e^(2t) + 3e^(-4t)

OpenStudy (turingtest):

yes :)

OpenStudy (anonymous):

Thank you so much for your time and patience!

OpenStudy (anonymous):

yeah I had model solutions but they were crappy :)

OpenStudy (turingtest):

you're welcome try again to read the link I gave you, it has great explanations and example problems note that there are many more constants of integration that poped up along the way, but they can be combined at the end if you know the situation you are dealing with well enough, so I ignored them in general though, remembering to add the constant of integration for every indefinite integral is indispensable in differential equations